OpenStudy (anonymous):

It takes a force of 78.0 N to compress the spring of a toy popgun 0.201 m to "load" a 0.145 kg ball. With what speed will the ball leave the gun? can anyone help me please

OpenStudy (anonymous):

First assume that all of the potential energy of the spring is transformed into kinetic energy of the ball.

OpenStudy (anonymous):

yeah but i just dont know how to do it at all im completely lost i just need the answer

OpenStudy (anonymous):

Calculate the amount of kinetic energy the ball receives.

OpenStudy (anonymous):

cant you give me the answer

OpenStudy (anonymous):

please

OpenStudy (anonymous):

W = Fs = E E = .5mv^2 <=> v = sqrt(E/.5m) v = sqrt(Fs/.5m) v = sqrt(78.0N*.201m/(.5*.145)) = 14.7053825 m / s approximately 15m/s

OpenStudy (anonymous):

Actually 14.7 m/s is fine.

OpenStudy (anonymous):

oh ok let me put it online

OpenStudy (anonymous):

no it did not work

OpenStudy (anonymous):

i plugged it my homework online

OpenStudy (anonymous):

It should be the correct answer. Try changing the rounding.

OpenStudy (anonymous):

Oh wait, is this a maths or a physics problem?

OpenStudy (anonymous):

physics

OpenStudy (anonymous):

in that case i don't see what's wrong with it.

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

can u solve another problem like that one just with different numbers

OpenStudy (anonymous):

to see if it works online

OpenStudy (anonymous):

this should be extremely simple. see if someone else gets a different result.

OpenStudy (anonymous):

no one has tried it only you

OpenStudy (anonymous):

hello help plz

OpenStudy (anonymous):

is the gun vertical or horizontal? this matters. Lets assume gun is being fired horizontally PE_i + KE_i = PE_f + KE_f PE_i = 1/2 k x^2 KE_i = 0 ball is not moving PE_f = 0 spring is not being compressed KE_f = 1/2 m v^2

OpenStudy (anonymous):

therefore PE_i = KE_f

OpenStudy (anonymous):

plz the answer

OpenStudy (anonymous):

v = sqrt(Fx/m) v = 10.398275719129325085729820157569 answer is wrong if being fired vertically

OpenStudy (anonymous):

wait the one you gave me now is wrong ???

OpenStudy (anonymous):

Isn't it v = sqrt(Fx/.5m) Because E=.5mv^2 <=> v = sqrt(E/.5m) ?

OpenStudy (anonymous):

thank you so much the duke you are right

OpenStudy (anonymous):

no answer depend on how the item is fired

OpenStudy (anonymous):

i was trying all night to get it thank you so much duke i truly appreciate it

OpenStudy (anonymous):

Zuige you mistake is the energy is not constant for a spring the spring energy is 1/2 k x^2 so the work is not simply force times displacement to get energy

OpenStudy (anonymous):

oh, of course. sorry for the mistake!

OpenStudy (anonymous):

thank u so much to all